An electron moving along the x axis has a position given by x = 16 te-t m, where t is in seconds. How far is the electron from the origin when it momentarily stops?
Answer #1

x = 16 te^-t
v = dx/dt = 16e^-t - 16te^-t = 0 when t = 1

x = 16 e^-1 = 5.88 m
Answer #2
x = 16 te-t m,
velocity=dx/dt=16*exp(-t)-16*t*exp(-t)
so v=0
at t=1
distance at that time=16*1*exp(-1)=5.886 m
Answer #3
t = 1

16/e = 5.897 m